OpenStudy (destinyyyy):

For f(x)=3x and g(x)=x+6, find the following functions. A. (f*g)(x) B. (g*f)(x) C. (f*g)(2) D. (g*f)(2)

OpenStudy (destinyyyy):

For A. I got 3x+18 For B. I got 3x+6 I'm stuck on C.

OpenStudy (james1769):

6+12

OpenStudy (destinyyyy):

I assume that's the answer for C.. I want to know how to solve it.

OpenStudy (destinyyyy):

@Nnesha

OpenStudy (james1769):

the same way u solved for the others

OpenStudy (destinyyyy):

If I solve it the same way I did the other two I get stuck at 3(g(2))

OpenStudy (james1769):

oh

Nnesha (nnesha):

substitute x for g(2) into f(x) function

OpenStudy (destinyyyy):

So 3(g(x)) ?

Nnesha (nnesha):

btw A is wrong

OpenStudy (destinyyyy):

How is it wrong? I followed the example.

Nnesha (nnesha):

when we multiply same bases we should `ADD` their exponents so 3x times x =?

OpenStudy (destinyyyy):

This is my first time doing this type of problem so I dont really know what im doing.

Nnesha (nnesha):

x is same as x to the one power apply the rule when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (destinyyyy):

Where are you getting a second x from?

Nnesha (nnesha):

we should multiply g(x) by f(x)

OpenStudy (destinyyyy):

My example told me to do this- (f*g)(x)= f(g(x))= 3(g(x))= 3(x+6)= 3x+18 It doesn't have x^2

Nnesha (nnesha):

well (f*g)(x) means f(x) times g(x)

Nnesha (nnesha):

where did you get that example ?

OpenStudy (destinyyyy):

My examples says (f*g)(x) means f(g(x)). My online class on Pearson. Its a system example.

Nnesha (nnesha):

f(g(x)) composition means (f o g)(x) where you should substitute x for g(x) into f(x) function

OpenStudy (destinyyyy):

Yeah.

Nnesha (nnesha):

so (f *g)(x) means f(x) times g(x) ^^^product of two function

OpenStudy (destinyyyy):

I entered my answer for the assignment and it says its correct.

Nnesha (nnesha):

what do you mean by `*` sign is it multiplication ?

OpenStudy (destinyyyy):

My examples for C does this- f(x)=3x g(x)=x+7 (f*g)(5), find the value of (f*g)(x) when x=5. (f*g)(x)= 3x+21 (f*g)(5)= 3(5)+21 36

OpenStudy (destinyyyy):

Yes * means multiply

Nnesha (nnesha):

looks at this http://home.windstream.net/okrebs/page42.html (f*g)(x) means f(x) times g(x)

OpenStudy (destinyyyy):

For that example above its not clear where the 21 comes from.

Nnesha (nnesha):

x+6 or x+7 ? g(x)^^

OpenStudy (destinyyyy):

x+7 is the example from the system. x+6 is the problem im working on. And the link you posted- I worked on that 2 problems ago.

Nnesha (nnesha):

well in that example the replaced x from with x+7 in f(x) function \[\large\rm f(x)=3\color{Red}{x} ~~~g(x) ~~~\large\rm f(x+7)=3(\color{ReD}{x+7})\]=\color{Red}{x+7}\]

Nnesha (nnesha):

so when you distribute parentheses by 3 you will get 3x +21 3 times 7 = 21

OpenStudy (destinyyyy):

Okay. I thought it was that but wasn't sure. So the answer for C is 33.

OpenStudy (destinyyyy):

Sorry 24..

Nnesha (nnesha):

yes 24 is correct according to that example

OpenStudy (destinyyyy):

And the answer for D. is 13

OpenStudy (destinyyyy):

Sorry 12.

Nnesha (nnesha):

yes right

OpenStudy (destinyyyy):

Okay. Thank you.

Nnesha (nnesha):

yw